Please help me solve this using matrices:
x + 2y - z = 2
2x - 2y + z = -1
6x +4y +3z = 5
Thank you so much, Lennie
: Please help me solve this using matrices:
x + 2y - z = 2
2x - 2y + z = -1
6x +4y +3z = 5
Thank you so much, Lennie

Answer by Fombitz(1744)

In Matrix notation, the problem is,
(<BR>
  matrix( 3, 3, <BR>
    1, 2, -1,<BR>
    2, -2, 1,<BR>
    6, 4, 3<BR>
  ))*<BR>
(<BR>
  matrix( 3, 1, <BR>
    x, <BR>
    y, <BR>
    z<BR>
  )<BR>
)=<BR>
(<BR>
  matrix( 3, 1, <BR>
    2, <BR>
    -1, <BR>
    5<BR>
  )<BR>
)
or simplified
[A][x]=[b]
[Ainv][A][x]=[Ainv][b]
[x]=[Ainv][b]
Find the inverse of matrix A, multiply by the right hand side to find the solution.
In this case,
[A]=(<BR>
  matrix( 3, 3, <BR>
    1, 2, -1,<BR>
    2, -2, 1,<BR>
    6, 4, 3<BR>
  ))
The inverse of [A] is,
[Ainv]=(1/30)*(<BR>
  matrix( 3, 3, <BR>
    10, 10, 0,<BR>
    0, -9, 3,<BR>
    -20, -8, 6<BR>
  ))
Then the matrix multiplication is,
[x]=(1/14)*(<BR>
  matrix( 3, 3, <BR>
    10, 10, 0,<BR>
    0, -9, 3,<BR>
    -20, -8, 6<BR>
  ))*(<BR>
  matrix( 3, 1, <BR>
    2, <BR>
    -1, <BR>
    5<BR>
  )<BR>
)
x=(1/30)(10*2+(10)(-1)+0(5))
x=(1/30)(20-10)
x=(1/30)(10)
x=1/3
y=(1/30)(0*2+(-9)(-1)+(3)(5))
y=(1/30)(9+15)
y=(1/30)(24)
y=4/5
z=(1/30)(20*2+(8)(-1)+(-6)(5))
z=(1/30)(40-8-30)
z=(1/30)(2)
z=-1/15
Your solution matrix is then
[x]=(<BR>
  matrix( 3, 1, <BR>
    1/3, <BR>
    4/5, <BR>
    -1/15<BR>
  )<BR>
)